OpenStudy (anonymous):

8.7 is 56% of what number

OpenStudy (phi):

8.7 is (means =) 56% of x (x is the name of "what number") 8.7 = 56/100 * x multiply both sides by 100/56 100*8.7/56 = x 870/56 = x can you finish ?

OpenStudy (anonymous):

15, thank you

OpenStudy (phi):

I would say 15.54 (rounded to the nearest 100th)

OpenStudy (phi):

or 15.5 rounded to the nearest tenth